sir, If sinA +cosA =2^0.5multiplied by sin(90-A);then determine cotA. thank you

sir,
If sinA +cosA =2^0.5multiplied by sin(90-A);then determine cotA.
thank you

Grade:10

1 Answers

SHAIK AASIF AHAMED
10 years ago
Hello student,
Please find the answer to your question below
Given sinA+cosA=\sqrt{2}sin(90-A)
So sinA+cosA=\sqrt{2}cosA
diving by cosA on both sides we get
TanA+1=\sqrt{2}
So tanA=\sqrt{2}-1
So cotA=(1/tanA)=(1/(\sqrt{2}-1))=\sqrt{2}+1
So cotA=\sqrt{2}+1